40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 13 12:21:09 2018, gautam, Update, ALS, PDFR laser checkout 
    Reply  Wed Jun 13 22:07:31 2018, gautam, Update, ALS, BeatMouth PDFR measurement IMG_7056.JPGBeatMouthPDFR.pdfBeatMouth_PDFRdata.tgz
       Reply  Fri Jun 15 14:22:05 2018, gautam, Update, ALS, BeatMouth PDFR measurement BeatMouthPDFR.pdfBeatMouth_PDFRdata.tgz
Message ID: 13973     Entry time: Fri Jun 15 14:22:05 2018     In reply to: 13957
Author: gautam 
Type: Update 
Category: ALS 
Subject: BeatMouth PDFR measurement 

I did the measurement with the BeatMouth open today. Main changes:

  • Directly pipe the RF output of the Menlo PDs to the Agilent, bypassing the 20dB coupler inside the BeatMouth.
  • Directly pipe the unused port of the Fiber Beamsplitter used to send light to the Menlo PD to an in-air collimator, which then sends the beam to the NF1611 reference detector.

So neglecting asymmetry in the branching ratio of the fiber beamsplitter, the asymmetry between the test PD optical path and the reference PD optical path is a single fiber mating sleeve in the former vs a collimator in the latter. In order to recover the expected number of 409 V/W for the Menlo PDs, we have to argue that the optical loss in the test PD path (fiber mating sleeve) are ~3x higher than in the NF1611 path (free space coupler). But at least the X and Y PDs show identical responses now. The error I made in the previously attached plot was that I was using the 20dB coupled output for the X PD measurement indecision.

Revised conclusion: The measured optoelectronic response of the Menlo PDs at 10s of MHz, of ~130 V/W, is completely consistent with the numbers I reported in this elog. So rogue polarization is no longer the culprit for the discrepancy between expected and measured RF beatnote power, it was just that the expectation, based on Menlo PD specs, were not accurate.#2 of the linked elog seems to be the most likely, although "broken" should actually be "not matching spec".


While killing time b/w measurements, I looked on the ITMY optical table and found that the NF1611 I mentioned in this elog still exists. It is fiber coupled. Could be a better substitute as a Reference PD for this particular measurement.

Quote:

I will repeat the measurement tomorrow by eliminating some un-necessary patch fiber cables, and also calibrating out the cable delays.

  • The setup shown in Attachment #1 was used because I didn't want to open up the BeatMouth.
  • But I can pipe the port of the BS not going to the FPD310 directly to the collimator, and that should reduce the systematic uncertainty w.r.t. power distribution between FPD310 and NF1611.
Attachment 1: BeatMouthPDFR.pdf  146 kB  Uploaded Fri Jun 15 15:29:48 2018  | Hide | Hide all
BeatMouthPDFR.pdf
Attachment 2: BeatMouth_PDFRdata.tgz  56 kB  Uploaded Fri Jun 15 15:30:00 2018
ELOG V3.1.3-